# Optimal Neural Compressors for the Rate-Distortion-Perception Tradeoff

### v1

@@ A.1 RCC Preliminaries
The one-shot RCC technique above enables immediate achievability results for informational quantities, such as the rate-distortion-perception function $R(D,P)$, by simulating the channel $P_{\hat{X}|X}$ that achieves the infimum in (1). Due to the fact that the reconstruction $\hat{X}_{K}$ has distribution equal to the channel $P_{\hat{X}|X}$, any constraint in the informational quantity, such as the expected distortion constraint, or the perception constraint, is automatically satisfied by RCC. The $I(X;\hat{X})$ term in (27) then becomes equal to the informational quantity $R(D,P)$. This is the approach used to prove achievability in Theis and Wagner [2021].
@@ A.2 RCC as Randomized VQ
The scheme in Def. 27 essentially chooses a random codebook at both the encoder and decoder. The encoder chooses a codeword $\hat{X}_{K}$ according to the criterion in (26), which may appear abstract at first glance. However, the following two propositions show that when the channel $P_{\hat{X}|X}$ used to simulate is chosen to be RD- or RDP-achieving, it becomes clear that (26) uses a minimum-distance codebook search similar to VQ.
@@
A closed-form for the density ratio of the RDP-achieving channel $P_{\hat{X}|X}$ for when the perception is measured by squared 2-Wasserstein (which is the focus of our paper) is currently not known. However, we conjecture that similar to (28) and (30), it will consist of a $e^{-\beta^{\prime}\operatorname{\Delta}(x,\hat{x})}$ term in the numerator, which will result in (26) having $\operatorname{\Delta}(x,\hat{X}_{i})$ in the objective. Another slight discrepancy to our setup is that the $n$-letter operational perception in Serra et al. [2023] for the $f$-divergence perception RDP function is measured in the weak-sense (see Sec. 2.2). We are focused on the strong-sense setting as it more faithfully describes practical usage.
@@ A.3 RCC Implementation Details
To simulate RCC, one can implement the scheme in Def. 27 and use (29) or (31) for finding the index to entropy-code. For the Gaussian source, the RDP-achieving channel and output marginal is given in Remark 4.2; we can directly implement (26) in closed-form since these distributions are Gaussian. This is what is done for the results in Fig. 9(a). Since it is not possible to generate an infinite number of samples $\hat{X}_{i}$, we generate a codebook $N=10,000$ samples instead. Following Li and El Gamal [2018], the index $K$ is entropy-coded using a Zipf distribution with parameter $\lambda=1+1/(I(X;\hat{X}+e^{-1}\log e+1)$. A full algorithm describing the encoding and decoding process can be found in Theis and Ahmed [2022] as well as Lei et al. [2022].
Appendix B Additional Empirical Results
@@
Figure 13: Comparing STE vs. noisy proxy for PD-NTC, on Speech.
Appendix C Proofs
C.1 Proof of Proposition 3.3
@@ Appendix C Proofs
$\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\bm{x}-\hat{\bm{x}}_{\mathsf{PD}}\|^{2}\right]=\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\bm{x}-Q_{\Lambda}(\bm{x})-\bm{u}\|^{2}\right]$ (33)
$=\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\bm{x}-Q_{\Lambda}(\bm{x})\|^{2}\right]+\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{u}\|^{2}]$ (34)
$=\sigma^{2}(\Lambda)+\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\bm{x}-Q_{\Lambda}(\bm{x})\|^{2}\right]$ (35)
@@
C.2 Proof of Theorem 4.4
Theorem C.1 (Optimality of SD-LTC for Gaussian sources (Thm. 4.4 in main text)).
@@
C.3 Proof of Theorem 4.6
We now consider the case when the encoder and decoder do not have access to infinite shared randomness. Unlike Thm. 4.4, Thm. 4.6 cannot make use of the additive channel equivalence, and we instead rely on results from lattice Gaussian coding [Ling and Belfiore, 2014]. We first introduce the concept of lattice Gaussians, then establish some results on the lattice covering radius as well as conditional moments of the chi-square distribution that will be used to prove Thm. 4.6.
Definition C.2 (Lattice Gaussian Distribution).
@@
For a more in-depth introduction, see Stephens-Davidowitz [2017]. The proof of Thm. 4.6 relies on known results regarding the lattice Gaussian second moment, entropy, and the fact that the sum of a lattice Gaussian and continuous Gaussian can be made arbitrarily close to a Gaussian in terms of total variation; see Ling and Belfiore [2014], Regev [2009], Banaszczyk [1993] for details.
@@
We now prove Thm. 4.6.
Theorem C.4 (PD-LTC achieves $R_{p}(D,0)$ for Gaussian sources (Thm. 4.6 in main text)).
Let $X_{1},X_{2},\dots\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma^{2})$. For any $D$ satisfying $0<D\leq 2\sigma^{2}$, there exists a sequence of PD-LTCs $\{(g_{a}^{(n)},g_{s}^{(n)},\Lambda^{(n)})\}_{n=1}^{\infty}$ such that the achieved rate, distortion, and perception satisfy
$\lim_{n\rightarrow\infty}\frac{1}{n}H(Q_{\Lambda^{(n)}}(g_{a}^{(n)}(X^{n})))=R_{p}(D,0),$ (75)
$\lim_{n\rightarrow\infty}\frac{1}{n}W_{2}^{2}(P_{X^{n}},P_{\hat{X}^{n}})=0,$ (77)
where $\hat{X}^{n}=g_{s}^{(n)}(Q_{\Lambda^{(n)}}(g_{a}^{(n)}(X^{n}))+\bm{u})$, $\bm{u}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda^{(n)}))$.
@@
We define the structure of the sequence of PD-LTCs as follows. Define the parameter $\nu:=\frac{D\sigma^{2}(4\sigma^{2}-D)}{D^{2}-4D\sigma^{2}+8\sigma^{4}}$. Let $g_{a}^{(n)}(\bm{v})=\alpha\bm{v}$, where $\alpha:=\frac{\sigma^{2}-\nu}{\sigma^{2}}$, and $g_{s}^{(n)}(\bm{v})=\beta\bm{v}$, where $\beta:=\frac{\sigma^{2}}{\sqrt{\sigma^{4}-\nu^{2}}}$. Choose $\Lambda^{(n)}$ to be a sequence of simultaneously AWGN-good and sphere-bound-achieving lattices [Ling et al., 2014, Zamir et al., 2014] such that the fundamental volume satisfes $V(\Lambda^{(n)})=\left(2\pi e\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}(1+\epsilon_{2})\right)^{\frac{n}{2}}$, where $\epsilon_{2}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}0$, following that of Ling and Belfiore [2014]. In the following, we denote the quantizers $Q_{\Lambda^{(n)}}$ as $Q$ with the dependence on the lattice implicit.
We first address the perception constraint. Fix $\epsilon>0$. Let $\tilde{\bm{x}}=\bm{y}+\bm{z}$, where $\bm{y}\sim\mathcal{N}_{\Lambda^{(n)}}(0,\sigma^{2}-\nu)$ is a lattice Gaussian distribution and $\bm{z}\sim\mathcal{N}(\bm{0},\nu\cdot I_{n})$. By triangle inequality,
$\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\hat{X}^{n}})$ $\leq\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\beta(Q(\alpha\tilde{\bm{x}})+\bm{u})})+\frac{1}{\sqrt{n}}W_{2}(P_{\hat{X}^{n}},P_{\beta(Q(\alpha\tilde{\bm{x}})+\bm{u})}).$ (78)
The second term on the right can be bounded by $\epsilon/2$ due to the following. From Ling and Belfiore [2014, Lemma 9], we know that $\delta_{\mathsf{TV}}(P_{\tilde{\bm{x}}},P_{\bm{x}})\leq\frac{\epsilon}{2}$ for $n$ sufficiently large, and thus $\delta_{\mathsf{TV}}\left(P_{\beta(Q(\alpha\tilde{\bm{x}})+\bm{u})},P_{\beta(Q(\alpha\bm{x})+\bm{u})}\right)\leq\frac{\epsilon}{2}$ as well, by data processing inequality. The result follows by the fact that convergence in $\delta_{\mathsf{TV}}$ implies convergence in $W_{2}$ [Arjovsky et al., 2017].
For the first term on the right, we again apply triangle inequality:
$\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\beta(Q(\alpha\tilde{\bm{x}})+\bm{u})})\leq\underbrace{\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\beta(\bm{y}+\bm{u})})}_{A}+\underbrace{\frac{1}{\sqrt{n}}W_{2}(P_{\beta(Q(\alpha\tilde{\bm{x}})+\bm{u})},P_{\beta(\bm{y}+\bm{u})})}_{B}.$ (79)
For term $A$, let $\tilde{\bm{z}}\sim\mathcal{N}\left(0,\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}I_{n}\right)$. note that
$\frac{1}{\beta\sqrt{n}}W_{2}(P_{X^{n}},P_{\beta(\bm{y}+\bm{u})})$ $=\frac{1}{\sqrt{n}}W_{2}\left(\mathcal{N}\left(0,\frac{\sigma^{4}-\nu^{2}}{\sigma^{2}}I_{n}\right),P_{\bm{y}+\bm{u}}\right)$ (80)
@@
$\leq\frac{1}{\sqrt{n}}W_{2}\left(\mathcal{N}\left(0,\frac{\sigma^{4}-\nu^{2}}{\sigma^{2}}I_{n}\right),P_{\bm{y}+\tilde{\bm{z}}}\right)+\frac{1}{\sqrt{n}}W_{2}\left(P_{\bm{y}+\tilde{\bm{z}}},P_{\bm{y}+\bm{u}}\right)$ (81)
$\leq\frac{\epsilon}{8\beta}+\frac{1}{\sqrt{n}}W_{2}(P_{\bm{u}},P_{\tilde{\bm{z}}})$ (82)

### v2

@@ A.1 RCC Preliminaries
Reverse channel coding (RCC), also known as channel simulation, devises schemes for sending a sample from a prescribed distribution with a rate constraint [Cuff, 2013, Li, 2024], and typically assumes shared randomness. Since RDP requires randomized reconstructions, RCC can be applied to RDP, and in fact is a technique used to prove RDP coding theorems; Theis and Wagner [2021] uses the one-shot RCC scheme of Li and El Gamal [2018], and Saldi et al. [2015], Wagner [2022] uses channel synthesis results of Cuff [2013]. We focus our discussion here on the particular RCC scheme of Li and El Gamal [2018], defined below.
@@
The one-shot RCC technique above enables immediate achievability results for informational quantities, such as the rate-distortion-perception function $R(D,P)$, by simulating the channel $P_{\hat{X}|X}$ that achieves the infimum in (1). Due to the fact that the reconstruction $\hat{X}_{K}$ has distribution equal to the channel $P_{\hat{X}|X}$, any constraint in the informational quantity, such as the expected distortion constraint, or the perception constraint, is automatically satisfied by RCC. The $I(X;\hat{X})$ term in (18) then becomes equal to the informational quantity $R(D,P)$. Applying this scheme to i.i.d. blocks yields the asymptotic result. This is the approach used to prove achievability in Theis and Wagner [2021].
@@ A.2 RCC as Randomized VQ
The scheme in Def. 18 essentially chooses a random codebook at both the encoder and decoder. The encoder chooses a codeword $\hat{X}_{K}$ according to the criterion in (17), which may appear abstract at first glance. However, the following two propositions show that when the channel $P_{\hat{X}|X}$ used to simulate is chosen to be RD- or RDP-achieving, it becomes clear that (17) uses a minimum-distance codebook search similar to VQ.
@@
Thus, when taking distortion into account, RCC can be seen as a sort of randomized VQ that finds the minimum-distance codeword (which are random). Despite its optimality for RDP, RCC is of high complexity (exponential in rate and dimension), and requires infinite shared randomness.
@@
A closed-form for the density ratio of the RDP-achieving channel $P_{\hat{X}|X}$ for when the perception is measured by squared 2-Wasserstein (which is the focus of our paper) is currently not known. However, we conjecture that similar to (19) and (21), it will consist of a $e^{-\beta^{\prime}\operatorname{\Delta}(x,\hat{x})}$ term in the numerator, which will result in (17) having $\operatorname{\Delta}(x,\hat{X}_{i})$ in the objective. Another slight discrepancy to our setup is that the $n$-letter operational perception in Serra et al. [2023] for the $f$-divergence perception RDP function is measured in the weak-sense (see Sec. 2). We are focused on the strong-sense setting as it more faithfully describes practical usage.
@@ A.3 RCC Implementation Details
To simulate RCC, one can implement the scheme in Def. 18 and use (20) or (22) for finding the index to entropy-code. For the Gaussian source, the RDP-achieving channel and output marginal is given in Remark 4.2; we can directly implement (17) in closed-form since these distributions are Gaussian. This is what is done for the results in Fig. 7(a). Since it is not possible to generate an infinite number of samples $\hat{X}_{i}$, we generate a codebook $N=10,000$ samples instead. Following Li and El Gamal [2018], the index $K$ is entropy-coded using a Zipf distribution with parameter $\lambda=1+1/(I(X;\hat{X}+e^{-1}\log e+1)$. A full algorithm describing the encoding and decoding process can be found in Theis and Ahmed [2022] as well as Lei et al. [2022].
Appendix B Additional Experimental Details
For the synthetic (i.i.d. Gaussian source), we set $n_{L}=8$, $g_{a}$ and $g_{s}$ to be linear functions, as the constructions in Thm. 4.3 and Thm. 4.5 suggest this to be sufficient for optimality, and use the $\mathbb{Z}_{8}$ and $E_{8}$ lattices. To cover the RDP tradeoff, we sweep a variety of $\lambda_{1},\lambda_{2}$ values. For the Speech and Physics datasets, we use MLPs for $g_{a}$ and $g_{s}$ of depth 3, hidden dimension 100, and softplus nonlinearities. For MNIST, we follow the same exact experimental setup of Blau and Michaeli [2019], including model architecture, and using the test Wasserstein distance via the discriminator neural network. For NTC models, we use the factorized $p_{\bm{y}}$ of Ballé et al. [2018], and for LTC models, we use the RealNVP normalizing flow [Dinh et al., 2017]. All models are trained for 100 epochs on the training data split, and reported metrics are averaged over the test split. Training is performed on a NVIDIA RTX5000 GPU.
Appendix C Additional Empirical Results
@@
Figure 8: RDP tradeoff of all models on MNIST. QSD-NTC/LTC-$\Gamma$ corresponds to QSD-NTC/LTC with a nesting ratio of $\Gamma$.
@@
Figure 12: Comparing straight-through estimator (STE) with hard quantization vs. noisy proxy for PD-NTC, on Speech.
Appendix D Proofs
D.1 Proof of Proposition 3.3
@@ Appendix D Proofs
$\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\bm{x}-\hat{\bm{x}}_{\mathsf{PD}}\|^{2}\right]=\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\bm{x}-Q_{\Lambda}(\bm{x})-s\bm{u}\|^{2}\right]$ (24)
$=\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\bm{x}-Q_{\Lambda}(\bm{x})\|^{2}\right]+\frac{1}{n}s^{2}\operatorname{\mathbb{E}}[\|\bm{u}\|^{2}]$ (25)
$=s^{2}\sigma^{2}(\Lambda)+\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\bm{x}-Q_{\Lambda}(\bm{x})\|^{2}\right]$ (26)
$\geq s^{2}\sigma^{2}(\Lambda)$ (27)
@@
D.2 Proof of Theorem 4.3
Theorem D.1 (Optimality of SD-LTC for Gaussian sources (Thm. 4.3 in main text)).
@@
D.3 Proof of Theorem 4.5
We now consider the case when the encoder and decoder do not have access to infinite shared randomness. Unlike Thm. 4.3, Thm. 4.5 cannot make use of the additive channel equivalence, and we instead rely on results from lattice Gaussian coding [Ling and Belfiore, 2014]. We first introduce several important concepts of lattice Gaussians, then establish several lemmas that will be used to prove Thm. 4.5.
Definition D.2 (Lattice Gaussian Distribution).
@@
For a more in-depth introduction, see Stephens-Davidowitz [2017]. The proof of Thm. 4.5 relies on known results regarding the lattice Gaussian second moment and entropy; see Ling and Belfiore [2014], Regev [2009], Banaszczyk [1993] for details.
Two tools which will be used throughout the proof are the (i) flatness factor of a lattice, and (ii) a vanishing error probability of maximum a posteriori (MAP) decoding of a lattice Gaussian signal sent over an AWGN channel. For (i), the flatness factor [Ling et al., 2014] of a lattice $\Lambda$ is defined as
$\epsilon_{\Lambda}(\gamma):=\max_{\bm{x}\in\mathcal{V}_{0}(\Lambda)}|V(\Lambda)\rho_{\gamma,\Lambda}(\bm{x})-1|,$ (61)
where $\rho_{\gamma,\Lambda}(\bm{x})=\frac{1}{(2\pi\gamma^{2})^{n/2}}\sum_{\bm{\lambda}\in\Lambda}e^{-\frac{\|\bm{x}-\bm{\lambda}\|^{2}}{2\gamma^{2}}}$. For (ii), suppose we wish to send $\bm{y}\sim\mathcal{N}_{\Lambda}(\bm{c},\sigma^{2}-\nu)$ over an AWGN channel with noise $\bm{z}\sim\mathcal{N}(0,\nu I_{n})$. Ling and Belfiore [2014] show that given the received signal $\tilde{\bm{x}}=\bm{y}+\bm{z}$, the MAP decoder of $\bm{y}$ given $\tilde{\bm{x}}$ is given by $Q_{\Lambda-\bm{c}}(\alpha\tilde{\bm{x}})$, where $\alpha=\frac{\sigma^{2}-\nu}{\sigma^{2}}$.
The high-level approach in the proof of Thm. 4.5 is to analyze the rate, distortion, and perception when compressing $\tilde{\bm{x}}=\bm{y}+\bm{z}$, following Liu et al. [2021, 2024]. This is because $P_{\tilde{\bm{x}}}$ approximates $\mathcal{N}(0,\sigma^{2}I_{n})$, with the approximation error given by the flatness factor $\epsilon_{\Lambda}\left(\sqrt{\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}}\right)$ [Ling and Belfiore, 2014, Regev, 2009]. With a vanishing MAP error probability, $Q_{\Lambda}(\alpha\tilde{\bm{x}})\approx\bm{y}$, and the overall system can be statistically analyzed using the properties of $\bm{y}$, which is a lattice Gaussian. We refer the reader to Ling et al. [2014], Ling and Belfiore [2014], Liu et al. [2021] for further details of lattice Gaussian coding.
@@
The next lemma bounds the second moment of $Q(\alpha\tilde{\bm{x}})$, which is used to bound the error terms.
Lemma D.4.
Let $\tilde{\bm{x}}=\bm{y}+\bm{z}$, $\bm{y}\sim\mathcal{N}_{\Lambda}(\bm{c},\sigma^{2}-\nu)$ and $\bm{z}\sim\mathcal{N}(0,\nu I_{n})$. Let $\alpha=\frac{\sigma^{2}-\nu}{\sigma^{2}}$. If $\{\Lambda^{(n)}\}_{n=1}^{\infty}$ is a sequence of lattices that is AWGN-good with vanishing error probability of MAP decoding of $\bm{y}$ given $\tilde{\bm{x}}$, then for any $\epsilon>0$,
$\operatorname{\mathbb{E}}[\|Q(\alpha\tilde{\bm{x}})\|^{2}]\leq\operatorname{\mathbb{E}}[\|\bm{y}\|^{2}]+\epsilon,$ (68)
for sufficiently large $n$.
@@
The second moment of $Q(\alpha\tilde{\bm{x}})$ satisfies
$\leq 8\|\bm{y}-\alpha\tilde{\bm{x}}\|^{4}+8\|\alpha\tilde{\bm{x}}\|^{4}$ (75)
where (74) is by triangle inequality and Cauchy-Schwarz, and (75) holds since $Q$ finds the closest vector to $\alpha\tilde{\bm{x}}$. We have that the expectations satisfy
$=O(n^{2}),$ (81)
and
$=O(n^{2}),$ (85)
for $n$ sufficiently large. This implies that
and therefore
@@
$\frac{1}{n}\sqrt{\operatorname{\mathbb{E}}\left[\|Q(\alpha\tilde{\bm{x}})\|^{4}\right]\operatorname{\mathbb{P}}(\mathcal{E})}<\epsilon$ (87)
for $n$ sufficiently large since $\lim_{n\rightarrow\infty}\operatorname{\mathbb{P}}(\mathcal{E})=0$. ∎
Before proving Thm. 4.5, we first prove a lemma which describes an achievable RDP region of pre/post-scaled lattice quantization with private dithering on the Gaussian source that makes the relationship between the flatness factors and the scaling $s$ explicit. Ensuring that all flatness factors vanish would imply a single-letter characterization of the achievable RDP region for the near-perfect perception regime, which is what Thm. 4.5 provides.
Lemma D.5.
Let $X_{1},X_{2},\dots\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma^{2})$. Let $\nu\in(0,\sigma^{2})$, $\alpha=\frac{\sigma^{2}-\nu}{\sigma^{2}}$, and $s>0$, $\beta>0$ be constants satisfying
$(\sigma^{2}-\nu)+s^{2}\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}=\frac{\sigma^{2}}{\beta^{2}}.$ (88)
Let $\tilde{\bm{x}}=\bm{y}+\bm{z}$, $\bm{y}\sim\mathcal{N}_{\Lambda}(\bm{c},\sigma^{2}-\nu)$ and $\bm{z}\sim\mathcal{N}(0,\nu I_{n})$. If $\{\Lambda^{(n)}\}_{n=1}^{\infty}$ is a sequence of lattices that is AWGN-good with vanishing error probability of MAP decoding of $\bm{y}$ given $\tilde{\bm{x}}$, and is also quantization-good, then for any $\epsilon>0$,
$\frac{1}{n}H(Q_{\Lambda^{(n)}}(\alpha X^{n}))\leq\frac{1}{1-4\epsilon_{z}}\left(\frac{1}{2}\log\frac{\sigma^{2}}{\nu}+\epsilon_{h}+\epsilon\right),$ (89)
$\frac{1}{n}\mathbb{E}[\|X^{n}-\hat{X}^{n}\|^{2}]\leq\frac{1}{1-4\epsilon_{z}}\left((1-\beta)^{2}(\sigma^{2}-\nu)+\nu+s^{2}\beta^{2}\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}+\epsilon\right),$ (90)
$\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\hat{X}^{n}})\leq\sqrt{8\sigma^{2}\epsilon_{u}}+\beta\sqrt{8\epsilon_{z}\left(2\frac{\sigma^{2}-\nu}{1-4\epsilon_{z}}+2s^{2}\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}\right)}+\epsilon,$ (91)
for $n$ sufficiently large, where $\hat{X}^{n}=\beta(Q_{\Lambda^{(n)}}(\alpha X^{n})+s\bm{u})$, $\bm{u}\sim\mathrm{Unif}(\mathcal{V}(\Lambda^{(n)}))$. In the above, $\epsilon_{h}=-\frac{\log(1-\epsilon_{1})}{n}+\frac{\pi(t^{-4}+1)\epsilon_{1}}{n(1-\epsilon_{1})}$, and
$\epsilon_{1}:=\epsilon_{\Lambda^{(n)}}\left(\frac{\sqrt{\sigma^{2}-\nu}}{\sqrt{\frac{\pi}{\pi-t}}}\right),\quad\epsilon_{z}:=\epsilon_{\Lambda^{(n)}}\left({\sqrt{\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}}}\right),\quad\epsilon_{u}:=\epsilon_{\Lambda^{(n)}}\left(\sqrt{\frac{(\sigma^{2}-\nu)\nu s^{2}}{\sigma^{2}+s^{2}\nu}}\right)$ (92)
are flatness factors of lattices $\Lambda^{(n)}$, with $0<t<1/e$.
Proof.
Since the lattice sequence is AWGN-good, choose the fundamental volume to satisfy $V(\Lambda^{(n)})=\left(2\pi e\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}(1+\epsilon_{2})\right)^{\frac{n}{2}}$, where $\epsilon_{2}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}0$, following that of Ling and Belfiore [2014]. In the following, we denote the quantizers $Q_{\Lambda^{(n)}}$ as $Q$ with the dependence on the lattice implicit.
Define :=$\mathcal{E}:=\{Q(\alpha\tilde{\bm{x}})\neq\bm{y}\}$ as the “error“ event of MAP decoding of $\bm{y}$ given $\tilde{\bm{x}}$ [Ling and Belfiore, 2014, Lemma 11]. By assumption, $\lim_{n\rightarrow\infty}\operatorname{\mathbb{P}}(\mathcal{E})=0$. Finally, since the lattices are sphere-bound-achieving, we have that the lattice NSM satisfies $\lim_{n\rightarrow\infty}G(\Lambda^{(n)})\rightarrow\frac{1}{2\pi e}$. The vanishing error probability and sphere-bound-achieving properties of the lattices will be used to establish the convergence.
We first address the perception constraint. By triangle inequality,
$\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\hat{X}^{n}})$ $\leq\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\beta(Q(\alpha\tilde{\bm{x}})+s\bm{u})})+\frac{1}{\sqrt{n}}W_{2}(P_{\hat{X}^{n}},P_{\beta(Q(\alpha\tilde{\bm{x}})+s\bm{u})}).$ (93)
We first bound the second term on the right. By Regev [2009, Claim 3.9] and Ling and Belfiore [2014, Lemma 9], $|dP_{\tilde{\bm{x}}}(\bm{x})-dP_{X^{n}}(\bm{x})|\leq 4\epsilon_{z}dP_{X^{n}}(\bm{x}),\forall\bm{x}$. By the change of variable formula, this implies $|dP_{\tilde{\alpha\bm{x}}}(\bm{x})-dP_{\alpha X^{n}}(\bm{x})|\leq 4\epsilon_{z}dP_{\alpha X^{n}}(\bm{x})$. Additionally, we have that for any $\bm{\lambda}\in\Lambda$,
[…]
and $\Pr\left(Q(\alpha\tilde{\bm{x}})=\bm{\lambda}\right)\geq(1-4\epsilon_{z})\Pr\left(Q(\alpha X^{n})=\bm{\lambda}\right)$ follows similarly. Thus
$|\Pr\left(Q(\alpha\tilde{\bm{x}})=\bm{\lambda}\right)-\Pr\left(Q(\alpha X^{n})=\bm{\lambda}\right)|\leq 4\epsilon_{z}\Pr\left(Q(\alpha X^{n})=\bm{\lambda}\right).$ (95)
Finally, we have
$dP_{Q(\alpha\tilde{\bm{x}})+s\bm{u}}(\bm{w})$ $=\sum_{\bm{\lambda}\in\Lambda}\Pr\left(Q(\alpha\tilde{\bm{x}})=\bm{\lambda}\right)dP_{s\bm{u}}(\bm{w}-\bm{\lambda})$ (96)
$\leq\sum_{\bm{\lambda}\in\Lambda}(1+4\epsilon_{z})\Pr\left(Q(\alpha X^{n})=\bm{\lambda}\right)dP_{s\bm{u}}(\bm{w}-\bm{\lambda})$ (97)
$=(1+4\epsilon_{z})dP_{Q(\alpha X^{n})+s\bm{u}}(\bm{w}),$ (98)
where the other direction follows similarly. Thus
$|dP_{Q(\alpha\tilde{\bm{x}})+s\bm{u}}(\bm{w})-dP_{Q(\alpha X^{n})+s\bm{u}}(\bm{w})|\leq 4\epsilon_{z}dP_{Q(\alpha X^{n})+s\bm{u}}(\bm{w}).$ (99)
for any $\bm{w}$. This implies
$\frac{1}{\beta\sqrt{n}}W_{2}\left(P_{\hat{X}^{n}},P_{\beta(Q(\alpha\tilde{\bm{x}})+s\bm{u})}\right)$ $=\frac{1}{\sqrt{n}}W_{2}\left(P_{Q(\alpha X^{n})+s\bm{u}},P_{Q(\alpha\tilde{\bm{x}}+s\bm{u})}\right)$ (100)
[…]
$\leq\sqrt{8\epsilon_{z}\frac{1}{n}\operatorname{\mathbb{E}}[\|Q(\alpha X^{n})+s\bm{u}\|^{2}]},$ (102)
where (101) is by Villani [2016, Thm. 6.15]. To bound the second norm, we first get
$\frac{1}{n}\operatorname{\mathbb{E}}[\|Q(\alpha X^{n})+s\bm{u}\|^{2}]$ $\leq 2\frac{1}{n}\operatorname{\mathbb{E}}[\|Q(\alpha X^{n})\|^{2}]+2\frac{1}{n}\operatorname{\mathbb{E}}[\|s\bm{u}\|^{2}]$ (103)
$=2\frac{1}{n}\operatorname{\mathbb{E}}[\|Q(\alpha X^{n})\|^{2}]+2s^{2}\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}.$ (104)
The second moment of $Q(\alpha X^{n})$ satisfies
$\operatorname{\mathbb{E}}[\|Q(\alpha X^{n})\|^{2}]$ $=\sum_{\bm{\lambda}}\|\bm{\lambda}\|^{2}\Pr(Q(\alpha X^{n})=\bm{\lambda})$ (105)
$=\sum_{\bm{\lambda}}\|\bm{\lambda}\|^{2}(\Pr\left(Q(\alpha\tilde{\bm{x}})=\bm{\lambda}\right)-\Pr\left(Q(\alpha X^{n})=\bm{\lambda}\right))+\sum_{\bm{\lambda}}\|\bm{\lambda}\|^{2}\Pr Q(\alpha\tilde{\bm{x}})=\bm{\lambda})$ (106)
$\leq 4\epsilon_{z}\operatorname{\mathbb{E}}[\|Q(\alpha X^{n})\|^{2}]+\operatorname{\mathbb{E}}[\|Q(\alpha\tilde{\bm{x}})\|^{2}],$ (107)
where the last step is by (95). Thus for $n$ sufficiently large,
$\frac{1}{n}\operatorname{\mathbb{E}}[\|Q(\alpha X^{n})\|^{2}]$ $\leq\frac{1}{1-4\epsilon_{z}}\frac{1}{n}\operatorname{\mathbb{E}}[\|Q(\alpha\tilde{\bm{x}})\|^{2}]$ (108)
$\leq\frac{1}{1-4\epsilon_{z}}(\sigma^{2}-\nu),$ (109)
where the last step is by Lemma D.4 and Banaszczyk [1993]. Combining with (102) and (104),